OpenStudy (anonymous):

What is the slope of the line between (-4, 4) and (-1, -2)

OpenStudy (anonymous):

can you put this in \[\frac{ y1-y2 }{ x1-x2 }\]

OpenStudy (anonymous):

either way works

OpenStudy (anonymous):

Go with @KlOwNlOvE otherwise you will get confused.. :)

OpenStudy (anonymous):

best to get as little negatives as possible

OpenStudy (anonymous):

It is beauty that you are able to take -ve as common from denominator too otherwise it could not work either way.. :P

OpenStudy (anonymous):

so -4--1 and --2

OpenStudy (anonymous):

Slope = 2

OpenStudy (anonymous):

2 or -2?

OpenStudy (anonymous):

-4-2 = ??

OpenStudy (eli_moses):

-2

OpenStudy (anonymous):

why are you doing -4-2? 4 looks positive to me

OpenStudy (anonymous):

-2 -4 = ??

OpenStudy (anonymous):

\[\frac{ 4-(-2) }{ -4-(-1) }\]

OpenStudy (anonymous):

That is for y2 - y1

OpenStudy (anonymous):

2 negatives will make a positive \[\frac{ 4+2 }{ -4+1 }\]

OpenStudy (anonymous):

\[m = \frac{-2 - 4}{-1 + 4}\]

OpenStudy (anonymous):

and thats for y2-y1^

OpenStudy (anonymous):

@KlOwNlOvE let me delete my replies, otherwise, user will get confused.. :)

OpenStudy (anonymous):

yall confused me a bit lol which way are we going

OpenStudy (anonymous):

You are doing y1 - y2..

OpenStudy (eli_moses):

??

OpenStudy (anonymous):

ok lets get started then

OpenStudy (anonymous):

\[\frac{ 4-(-2) }{ -4-(-1) }\]

OpenStudy (eli_moses):

@KlOwNlOvE your formula is wrong

OpenStudy (anonymous):

formula can go either way as long as x1 is lined up with y1

OpenStudy (eli_moses):

ok

OpenStudy (anonymous):

@Eli_Moses, we two are enough to make a user get confused easily, you won't want to join us, otherwise, user will not only close this question, he or she will delete his or her OS account too.. :P

OpenStudy (eli_moses):

lol ok

OpenStudy (anonymous):

\[\frac{ y1-y2 }{ x1-x2 }\frac{ y2-y1 }{ x2-x1 } but no \frac{ y2-y1 }{ x1-x2 }\]

OpenStudy (anonymous):

@KlOwNlOvE the formula you are using is giving right results but it is not standard one.. We generally chose final - initial, so y1 is initial coordinate and y2 is final..

OpenStudy (anonymous):

a guy thatll confuse someone is phi but really good teacher

OpenStudy (anonymous):

I can too but I don't.. Users here are already confused, so why to double their trouble..!! :P

OpenStudy (anonymous):

He was teaching me how to do weird calculus yesterday idk how many times i got confused

OpenStudy (eli_moses):

\[\frac{ y _{2}-y _{1} }{ x _{2} - x _{1} }\] \[(-4, 4) (-1, -2)\] \[(x _{1}, y _{1})(x _{2}, y _{2})\] so then \[\frac{ (-2) - 4 }{ (-1)-(-4) }\] or \[\frac{ (-2)-4 }{ (-1)+4 }\] since negative and negative equal positive. now solve... \[\frac{ (-2) - 4 }{ (-1) +4 } = \frac{ -6 }{ 3 } = \frac{ -2 }{ 1 } = -2\] so the answer is -2. Your welcome. :P

OpenStudy (anonymous):

Checked: Slope: k = dy/dx = -2 http://www.hackmath.net/en/calculator/line-slope?x0=-4&y0=4&x1=-1&y1=-2&submit=Solve